# From Chaos to Coherence: Effects of High-Order Synaptic Correlations on Neural Dynamics

### v1

@@ Abstract
Recurrent Neural Network models have elucidated the interplay between structure and dynamics in biological neural networks, particularly the emergence of irregular and rhythmic activities in cortex. However, most studies have focused on networks with random or simple connectivity structures. Experimental observations find that high-order cortical connectivity patterns affect the temporal patterns of network activity, but a theory that relates such complex structure to network dynamics has yet to be developed. Here, we show that third- and higher-order cyclic correlations in synaptic connectivities greatly impact neuronal dynamics. Specifically, strong cyclic correlations in a network suppress chaotic dynamics, and promote oscillatory or fixed activity. The change in dynamics is related to the form of the unstable eigenvalues of the random connectivity matrix. A phase transition from chaotic to fixed or oscillatory activity coincides with the development of a cusp at the leading edge of the eigenvalue support. We also relate the dimensions of activity to the network structure.
@@
A central goal of theoretical neuroscience is to relate the structure, dynamics, and function of biological neural networks [1, 2, 3, 4, 5, 6, 7, 8]. Model networks with random, Erdős–Rényi connectivity have been studied extensively, and exhibit a range of explainable dynamical behaviors, from fixed points and periodic orbits to highly chaotic states [1]. However, biological neuronal networks have highly structured connectivity [9, 10, 11, 12, 13, 14, 15, 16]. While much effort has been devoted to understanding how such structure impacts neural dynamics and function, most work has focused on the impact of pairwise correlations, or structures encompassing the entire network [13, 17, 18, 19, 20, 21].
Figure 1: Behavior of networks with third-order cyclic correlations: (a) Distributions of complex connectivity matrix eigenvalues $\lambda$ at different correlation strengths $\rho$ for effective gain $g_{\text{eff}}=1.25$. The red lines correspond to Re $\lambda=1$; eigenvalues to the right of the red line are unstable. (b) Activities of five typical neurons at each value of $\rho$. (c) Heat map showing probabilities of observing chaotic, oscillatory, or fixed point activity (red, green, and blue, correspondingly) as a function of $g_{\text{eff}}$ and $\rho$. Dashed lines correspond to $\rho=\pm 0.5$. (d) Averaged normalized attractor dimension (participation ratio) at constant values of $g_{\text{eff}}$ as a function of $\rho$.
@@
Here, using numerical simulations we show that third- and higher-order cyclic correlations in synaptic connectivities have a strong impact on neuronal dynamics. We then characterize the connectivity structure through eigenvalue spectra and relate it to the network dynamics, showing how the emergence of irregular activity is strongly affected by high-order correlations in connectivity. We find that strong, positive high-order correlations tend to stabilize neuronal dynamics and reduce the dimensionality of activity. In contrast, strong, negative high-order correlations lead to high-frequency rhythmic behavior. Our results suggest that networks with such connectivity structure do not display chaotic behavior, unlike unstructured [1] and partially symmetric networks [21].
@@
where $x_{i}$ is the membrane potential of neuron $i$, $\phi(x_{j})=\tanh(x_{j})$ is the activation function of neuron $j$, and $w_{ij}$ is the weight of the synaptic connection from neuron $j$ to neuron $i$. Weights satisfy $\langle w_{ij}\rangle=0,\langle w^{2}_{ij}\rangle=\frac{g^{2}}{N},$ and
@@
where the angular brackets delimited-⟨$\langle$..$\rangle$ … ⟩ denote the ensemble average. The gain parameter $g$ determines the coupling strength, $\rho$ controls the strength of high-order directed cyclic correlations, and $\alpha$ denotes correlation order. We used a modification of the algorithm proposed by Aceituno, et al. [36] to generate matrices with the desired statistics (see section I.1).
Figure 2: Distribution of random connectivity matrix eigenvalues when $g_{\text{eff}}=1.1$ for (a) $\rho\approx 0.23$, (b) $\rho\approx 0.45$, and (c) $\rho\approx 0.76$. Red lines at $\mbox{Re}\;\lambda=1$ mark limit of stability for eigenvalues. Insets show expanded views of the distributions of the unstable eigenvalues at the tips that control the network dynamics.
The eigenvalue spectrum of random matrices with high-order cyclic correlations obeys a Hypotrochoid Law [32, 33, 36]: The eigenvalue density when $\rho\ll 1$ and $N\rightarrow\infty$ has support inside the hypotrochoid $z(\phi)=g(e^{i\phi}+\rho e^{-i(\alpha-1)\phi})$. In contrast to uncorrelated or pairwise correlated weights, the eigenvalue density when $\alpha\geq 3$ is not uniform within the hypotrochoid,  [40, 41, 42, 36]. At finite values of $|\rho|$ we observe deviations from the Hypotrochoid Law (See Fig. S2): At small values of $|\rho|$, eigenvalues can be found just outside the hypotrochoid, and at $|\rho|>\rho_{\text{c}}$, where $\rho_{\text{c}}\equiv(\alpha-1)^{-1}$, hypotrochoids have loops, which the eigenvalue density does not follow.
The Hypotrochoid Law, nevertheless, provides an approximation of the real part of the leading (rightmost) eigenvalue, which we call the effective gain, $g_{\text{eff}}$. In general, the phase of the hypotrochoid that corresponds to the leading eigenvalue is a solution of $U_{\alpha-2}(\cos(\phi^{*}))=-(\rho(\alpha-1))^{-1}$, where $U_{\alpha}(x)$ are the Chebyshev polynomials of the second kind. For $\rho\geq\rho_{\text{f}}\equiv-(\alpha-1)^{-2}$, and $\phi^{*}=0$, while for $\rho<\rho_{\text{f}}$ the solution $\phi^{*}$ for the leading edge of the approximate eigenvalue support has a more complex form. Then $g_{\text{eff}}$ is the real part of the support at $\phi^{*}$. For $\alpha=3$,
@@
In contrast to networks with uncorrelated weights, we find that high-order cyclic correlations ($\alpha\geq 3$) suppress the emergence of chaotic activity near the onset of nontrivial dynamics. Fig. 1 Shows results for networks with $\alpha=3$ and $N=1600$ neurons. Figs. 1(a)–(b) show representative network activity near this onset. When weights are uncorrelated, $\rho=0$, chaotic activity emerges near the onset of instability of the origin, and persists when correlations are weak and positive, $\rho\approx 0.23$. However, solutions typically converge to a fixed point when correlations are strong ($\rho\approx 0.76$ in Fig. 1(b)). Similarly, chaotic activity persists when correlations are weak and negative, $\rho\approx-0.23$, but is replaced by high frequency oscillations when correlations are strong and negative, $\rho\approx-0.76$.
@@
Figure 1(c) shows the probability of observing chaotic, oscillatory, or fixed point dynamics as $g_{\text{eff}}$ and $\rho$ are varied with the intensity of the different colors representing the empirical probability of finding each state. To obtain these probabilities we smoothed the results from 300 simulations at each point of a parameter grid with spacings $\Delta g_{\text{eff}}\approx 0.29$ and $\Delta\rho\approx 0.08$, with each simulation using a different network realization (See section I.2 for details.) The fact that oscillations or fixed points are sometimes found near the onset of instability of the quiescent fixed point when $|\rho|<\rho_{\text{c}}$ is likely a finite size effect. The range of $g_{\text{eff}}$ for which we find fixed points or oscillations after this onset decreases in width as $1/\sqrt{N}$ for $\rho=0$ [1]. We find that this range also decreases with $N$ for $|\rho|<\rho_{\text{c}}$. (See Fig. S3(a)-S3(b).)
@@
In recurrent neural networks, the dimensionality of attractors is constrained by the network architecture [19]. A common measure of dimensionality is the participation ratio normalized by network size [43, 44, 19]. Fig. 1(d) shows that this normalized participation dimension peaks at $\rho\approx-0.3$ for different values of $g_{\text{eff}}$, and decreases considerably from this peak at larger or smaller values of weight correlations, $\rho$. Thus, even when solutions are chaotic, third order cyclic correlations strongly impact network dynamics.
@@
What causes the observed stabilization of network dynamics? We hypothesize that it is due to the distribution of unstable eigenvalues, i.e. eigenvalues $\lambda$ with $\mbox{Re}\;\lambda>1$. The shape of the support of the eigenvalue spectrum, which is approximately given by the hypotrochoid curve, changes with $\rho$. Fig. 1(a) shows examples for $\alpha=3$. In general, the support has $\alpha$ “vertices”. The spectral distributions at $\rho$ and $-\rho$ are identical up to a rotation by $\pi/\alpha$. For $\rho>\rho_{\text{f}}$, one of the vertices is aligned with the real axis, while for $\rho<\rho_{\text{f}}$, the real axis is half way between two vertices. When $|\rho|<\rho_{\text{c}}$, the hypotrochoid is smooth at the vertices, but when $|\rho|\geq\rho_{\text{c}}$, the hypotrochoid has cusps. This change in the morphology results in a phase transition in the dynamics at the onset of instability of the quiescent fixed point.
@@
Figure 2 shows how the nonuniform eigenvalue density within the support changes as a function of $\rho$ for $\alpha=3$ ($100$ realizations, $N=6400$). Eigenvalues accumulate near the tips of the vertices as $|\rho|$ increases. For all $\rho$ there is an excess of real eigenvalues. For networks with pairwise correlations ($\alpha=2$), this excess density is known to decay as $\sim 1/\sqrt{N}$ [2]. For weak correlations, the eigenvalue distribution is nearly uniform, e.g., at $\rho\approx 0.23$ as shown in Fig. 2(a). As $\rho$ increases, the eigenvalues accumulate near the tips of the vertices, e.g., at $\rho\approx 0.45$ as shown in Fig. 2(b). However, for $\rho>\rho_{\text{c}}=0.5$, when the vertices are cusps, the eigenvalue density is maximal at a point in the middle of each cusp, away from the tips. Here, the eigenvalue density decays slowly towards the tips from the maxima, as can be seen for $\rho\approx 0.76$ in Fig. 2(c).
@@
The leading eigenvalues in the spectrum control the stability of the dynamics. For $\rho=0$, the spectrum is given by the circular law [1, 45], and the leading eigenvalues are at the leading edge of the circle. In the limit of large $N$, many eigenvalues become unstable simultaneously at the onset of instability of the quiescent fixed point where $g_{\text{eff}}\approx 1^{+}$, leading to chaos  [1]. The same appears to be true for all $|\rho|<\rho_{\text{c}}$: many eigenvalues become unstable simultaneously and chaotic dynamics emerge with high probability. For $\rho\approx\rho_{\text{f}}$, where the leading edge of the spectrum is flat (vertical), dynamics has maximal dimensionality (See Fig. 1(d)). At this value of $\rho$, the probability of observing chaos when the origin becomes unstable is also maximal.
@@
For $\rho>\rho_{\text{f}}$, the vertex at the leading edge of the support is along the real axis. As $\rho$ increases, the vertex becomes increasingly pointed, the eigenvalues accumulate around the real axis, the magnitudes of their imaginary parts decrease, and the fluctuations in the dynamics slow. (See insets of Figs. 2(a) and 2(b).) For $\rho<\rho_{\text{f}}$, the leading edge of the spectrum occurs at two symmetric vertices on either side of the real axis. Here, as $\rho$ decreases, the eigenvalues accumulate near the two vertices. In both cases, for $|\rho|>\rho_{\text{c}}$, when there are cusps at the tips of the vertices, eigenvalues do not accumulate at the edge of the support. Instead, the eigenvalue density decays toward the tips of the cusps. Because of this, the number of unstable eigenvalues increases slowly with $g_{\text{eff}}$ after the onset of instability of the quiescent fixed point, even in the limit of large $N$. (See inset of Fig. 2(c).) This slow increase corresponds with the high probability of observing fixed points or oscillations in the dynamics. Thus, at $\rho=\pm\rho_{\text{c}}$, where there is a morphological change at the leading edge of the eigenvalue support, we conjecture that there are critical points where phase transitions in the dynamics at onset occur. This conclusion is consistent with what is known about networks with pairwise correlations ($\alpha=2$), where chaos occurs at onset for all $|\rho|<1$, but fixed point behavior occurs at $\rho=\rho_{\text{c}}=1$ [21, 46].
@@
Although chaotic dynamics do not occur at the onset of instability of the origin for strong correlations where $|\rho|>\rho_{\text{c}}$, it does appear when $g_{\text{eff}}$ is large, e.g., for $|\rho|\approx 0.75$ chaos is still observed with high probability when $g_{\text{eff}}\gtrsim 3$ in Fig. 1(c). The value of $g_{\text{eff}}$ for which chaos is first observed decreases as $|\rho|\rightarrow\rho_{\text{c}}$, presumably to $\sim 1^{+}$ in the limit of large $N$. However, simulations are complicated in this region because the transient time increases with both $g_{\text{eff}}$ and $N$, and it becomes difficult to distinguish between chaotic behavior, and fixed point or oscillatory dynamics with long transients.
@@
Numerical investigations of networks with $\alpha=4,5\ \text{and}\ 6$ confirm the existence of phase transitions in the dynamics at onset and critical points at $\rho=\pm\rho_{\text{c}}$ consistent with the above predictions. However, at values of $|\rho|\gg\rho_{\text{c}}$, deviations from the Hypotrochoid Law become increasingly large with $\alpha$. For these networks, at such high values of $|\rho|$, our approximation for the gain at the onset of nontrivial dynamics, $g_{\text{eff}}$, becomes inaccurate.
The network dynamics within the region where chaos is observed are strongly influenced by high-order correlations. Fig. 3 shows the average attractor dimensionality and Lyapunov exponents $\lambda_{\text{LE}}$ as a function of $\alpha$, at fixed values of $|\rho|\approx 0.18$ and $g_{\text{eff}}$ The data are averages over $50$ realizations of networks with $N=1600$. (See Fig. S4 for the corresponding representative eigenvalue distributions.) Both quantities behave similarly for both positive and negative correlations. For the positive value of $\rho$, both quantities decrease monotonically with $\alpha$. For the negative value of $\rho$, the quantities peak near $\alpha=3$. For $\alpha\leq 3$, the correlation strength is greater than and approaching $\rho_{\text{f}}$, while for $\alpha>3$, it is less than and diverging from $\rho_{\text{f}}$. These results are consistent with Fig. 1(d) that shows the largest dimensionality occurs at $\rho\approx\rho_{\text{f}}$, where the leading edge of the support is flat and many eigenvalues become unstable simultaneously at onset.
@@
Figure 4: Phase diagram of dynamics at the onset of the instability of the origin as a function of correlation strength $\rho$ and correlation order $\alpha$ in the limit of $N\rightarrow\infty$. Solid lines correspond to ±$\pm\rho_{\text{c}}$, where the eigenvalue support of the synaptic connectivity matrix develops cusps. Dashed line corresponds to $\rho_{\text{f}}$, where the leading edge of the support is flat and the dimensionality of the dynamics is maximal.
@@
Thus, we conjecture the phase diagram in Fig. 4 for the dynamics at the onset of instability of the quiescent fixed point as a function of $\rho$ and $\alpha$, where $\alpha$ is extended to a real variable. Lines of critical points (solid lines) separate the chaotic phase (red) from the phases where fixed points (blue) and stable oscillations (green) occur. The dimensionality of the dynamics and the Lyapunov exponents are maximal along a line in the midst of the chaotic region (dashed line).

### v2

@@ Abstract
Recurrent Neural Network models have elucidated the interplay between structure and dynamics in biological neural networks, particularly the emergence of irregular and rhythmic activities in cortex. However, most studies have focused on networks with random or simple connectivity structures. Experimental observations find that high-order cortical connectivity patterns affect the temporal patterns of network activity, but a theory that relates such complex structure to network dynamics has yet to be developed. Here, we show that third- and higher-order cyclic correlations in synaptic connectivities greatly impact neuronal dynamics. Specifically, strong cyclic correlations in a network suppress chaotic dynamics and promote oscillatory or fixed activity. This change in dynamics is related to the form of the unstable eigenvalues of the random connectivity matrix. A phase transition from chaotic to fixed or oscillatory activity coincides with the development of a cusp at the leading edge of the eigenvalue support. We also relate the dimensions of activity to the network structure.
@@
A central goal of theoretical neuroscience is to relate the structure, dynamics, and function of biological neural networks [1, 2, 3, 4, 5, 6, 7, 8]. Model networks with random, Erdős–Rényi connectivity have been studied extensively, and exhibit a range of dynamical behaviors, from fixed points and periodic orbits to highly chaotic states [1]. However, biological neuronal networks have highly structured connectivity [9, 10, 11, 12, 13, 14, 15, 16]. While much effort has been devoted to understanding how such structure impacts neural dynamics and function, most work has focused on the impact of pairwise correlations, or structures encompassing the entire network [13, 17, 18, 19, 20, 21, 22].
Figure 1: Dynamics of networks with third-order cyclic correlations $(\alpha=3)$: (a) Distributions of connectivity matrix eigenvalues, $\lambda,$ at different correlation strengths, $\rho,$ for effective gain $g_{\text{eff}}=1.25$. The red lines correspond to Re $\lambda=1$; eigenvalues to the right of the red line are unstable. (b) Activities of four typical neurons at each value of $\rho$. (c) Heat map showing probabilities of observing chaotic, oscillatory, or fixed point activity (red, green, and blue, respectively) as a function of $g_{\text{eff}}$ and $\rho$. White dashed lines correspond to $\rho=\pm 0.5$. The black dotted line at $\rho\approx-0.28$ depicts the value at which the probability of observing chaotic behavior was highest when $g_{\text{eff}}=1.2$. (d) Averaged normalized attractor dimension (participation ratio) at constant values of $g_{\text{eff}}$ as a function of $\rho$.
@@
Here, using numerical simulations we show that third- and higher-order cyclic correlations in synaptic connectivities have a strong impact on neuronal dynamics. We then characterize the connectivity structure through its eigenvalue spectra and relate it to the network dynamics. We find that the emergence of irregular activity is strongly affected by high-order correlations in connectivity: Strong, positive high-order correlations tend to stabilize neuronal dynamics and reduce the dimensionality of activity. In contrast, strong, negative high-order correlations lead to high-frequency rhythmic behavior. Our results suggest that networks with such connectivity structure do not display chaotic behavior, unlike unstructured [1] and partially symmetric networks [22].
@@
where $x_{i}$ is the membrane potential of neuron $i$, $\phi(x_{j})=\tanh(x_{j})$ is the activation function of neuron $j$, and $w_{ij}$ is the weight of the synaptic connection from neuron $j$ to neuron $i$. Weights satisfy $\langle w_{ij}\rangle=0,\langle w^{2}_{ij}\rangle=g^{2}/N,$ and
@@
where the angular brackets $\langle$$\rangle$ denote the ensemble average. The gain parameter $g$ determines the coupling strength, $\rho$ controls the strength of high-order directed cyclic correlations, and $\alpha$ denotes correlation order. We used a modification of the algorithm proposed by Aceituno, et al. [37] to generate matrices with the desired statistics (see Sec.˜I.1 of supplementary materials (SM) for details).
Figure 2: Distribution of connectivity matrix eigenvalues when $\alpha=3,g_{\text{eff}}=1.1,$ and (a) $\rho\approx 0.23$, (b) $\rho\approx 0.45$, and (c) $\rho\approx 0.76$. Red lines at $\mbox{Re}\;\lambda=1$ mark the limit of stability for eigenvalues. Insets show expanded views of the distributions of the unstable eigenvalues at the tips that control the network dynamics.
The eigenvalue spectrum of random matrices with high-order cyclic correlations obeys a Hypotrochoid Law [33, 34, 37]: The eigenvalue density when $|\rho|\ll 1$ and $N\rightarrow\infty$ has support inside the hypotrochoid $z(\phi)=g(e^{i\phi}+\rho e^{-i(\alpha-1)\phi})$. In contrast to uncorrelated or pairwise correlated weights, the eigenvalue density when $\alpha\geq 3$ is not uniform within the hypotrochoid,  [41, 42, 43, 37]. At finite values of $\rho$, we observe deviations from the Hypotrochoid Law (see Fig. S2 of SM): At small values of $|\rho|$, eigenvalues can be found just outside the hypotrochoid, and at $|\rho|>\rho_{\text{c}}$, where $\rho_{\text{c}}\equiv(\alpha-1)^{-1}$, hypotrochoids have loops, which the eigenvalue density does not follow.
The Hypotrochoid Law, nevertheless, provides an approximation of the real part of the leading (rightmost) eigenvalue, which we call the effective gain, $g_{\text{eff}}$. In general, the phase of the hypotrochoid that corresponds to the leading eigenvalue is a solution of $U_{\alpha-2}(\cos(\phi^{*}))=-(\rho(\alpha-1))^{-1}$, where $U_{\alpha}(x)$ are the Chebyshev polynomials of the second kind. At
$\rho_{\text{f}}\equiv-(\alpha-1)^{-2},$ (3)
the leading edge of the eigenvalue distribution is on the real axis, and the curvature of the support at that point is zero. For $\rho\geq\rho_{\text{f}}$, the leading edge of the eigenvalue support has a unique solution corresponding to $\phi^{*}=0$, while for $\rho<\rho_{\text{f}}$ it has two solutions. The effective gain, $g_{\text{eff}},$ is the real part of the support at $\phi^{*}$, and for $\alpha=3$,
@@
In contrast to networks with uncorrelated weights, we find that high-order cyclic correlations ($\alpha\geq 3$) suppress the emergence of chaotic activity near the onset of nontrivial dynamics. Figs. 1(a)–(b) show representative network activity near this onset. When weights are uncorrelated, $\rho=0$, chaotic activity emerges near the onset of instability of the origin, and persists when correlations are weak and positive. However, solutions typically converge to a fixed point when correlations are strong ($\rho\approx 0.76$ in Fig. 1(b)). Similarly, chaotic activity persists when correlations are weak and negative, $\rho\approx-0.23$, but is replaced by high frequency oscillations when correlations are strong and negative, $\rho\approx-0.76$.
@@
Figure 1(c) shows the probability of observing chaotic, oscillatory, or fixed point dynamics as $g_{\text{eff}}$ and $\rho$ are varied with the intensity of the different colors representing the empirically determined probability of each state. To obtain these probabilities we smoothed the results from 300 simulations at each point of a parameter grid with spacings $\Delta g_{\text{eff}}\approx 0.29$ and $\Delta\rho\approx 0.08$, with each simulation using a different network realization (see Sec.˜I.2 of SM for details). The fact that oscillations or fixed points are sometimes found near the onset of instability of the quiescent fixed point when $|\rho|<\rho_{\text{c}}$ is presumably a finite size effect, as is known analytically to be the case for $\rho=0$ [1]. In Fig. 1, the results are shown for $N=1600$. The range of $g_{\text{eff}}$ for which fixed points or oscillations occur after this onset decreases in width as $1/\sqrt{N}$ for $\rho=0$ [1]. We find that this range also decreases with $N$ for $|\rho|<\rho_{\text{c}}$ (see Fig. S3(a)-S3(b) of SM). Numerical scaling shows that non-chaotic behavior vanishes in the limit of $N\rightarrow\infty$ as $g_{\mathrm{eff}}\rightarrow 1$ and $|\rho|<\rho_{\text{c}}$ (see Fig. S4(a)-S4(b) of SM). In contrast, when $|\rho|>\rho_{\text{c}}$, near the onset we still frequently observed non-chaotic solutions numerically even when $N$ was large (see Fig. S3(c) of SM). However, even for $|\rho|>\rho_{\text{c}}$, chaotic solutions were detected far from the onset, at large values of $g_{\mathrm{eff}}>1$ (see Fig. S4(c) of SM). Although in this regime, transient effects can cause an overestimation of the probabilities of observing chaotic dynamics (see Sec.˜I.3 of SM).
@@
In recurrent neural networks, the dimensionality of attractors is constrained by the network architecture [20]. The dimensions of the sub-space spanned by these attractors can affect a network’s ability to learn [44], and reflect properties of external stimuli [45]. A common measure of dimensionality is the participation ratio normalized by network size [46, 47, 20]. Fig. 1(d) shows that this normalized participation dimension peaks at $\rho\approx-0.3$ for different values of $g_{\text{eff}}$, and decreases considerably from this peak at other values of $\rho$. Thus, even when solutions are chaotic, third order cyclic correlations strongly impact network dynamics.
@@
What causes the observed stabilization of network dynamics? We hypothesize that this is due to the distribution of unstable eigenvalues, i.e. eigenvalues $\lambda$ with $\mbox{Re}\;\lambda>1$. The shape of the support of the eigenvalue spectrum, approximately described by the hypotrochoid curve, changes with $\rho$ (See Fig. 1(a) for $\alpha=3$). In general, the support has $\alpha$ “vertices”. The eigenvalue distributions at $\rho$ and $\rho$ are identical up to a rotation by $\pi/\alpha$. For $\rho>\rho_{\text{f}}$ (see Eq.˜3), the leading edge is on the real axis, while for $\rho<\rho_{\text{f}}$, the leading edge occurs at two points with the real axis half way between them. When $|\rho|<\rho_{\text{c}}$, the eigenvalue support is smooth at the vertices, but when $|\rho|\geq\rho_{\text{c}}$, the support has cusps. We observed numerically that this change in the morphology coincides with a phase transition in the dynamics at the onset of instability of the quiescent fixed point. Thus, $\rho=\pm\rho_{\text{c}}$ are critical points.
@@
This morphological change as a function $\rho$ also involves the development of a nonuniform eigenvalue density as shown in Fig. 2, which is calculated for $\alpha=3$ ($100$ realizations, $N=6400$). In particular, eigenvalues accumulate near the tips of the vertices as $|\rho|$ increases, with an excess of real eigenvalues observed for all values of $\rho$. For networks with pairwise correlations ($\alpha=2$), this excess of real values is known to decay as $\sim 1/\sqrt{N}$, and thus disappears as the network grows in size [2]. For weak correlations, the eigenvalue distribution is nearly uniform, e.g., at $\rho\approx 0.23$ as shown in Fig. 2(a). As $\rho$ increases, the eigenvalues accumulate near the tips of the vertices, e.g., at $\rho\approx 0.45$ as shown in Fig. 2(b). However, for $\rho>\rho_{\text{c}}=0.5$, when the vertices are cusps, the eigenvalue density is maximal at a point in the middle of each cusp, away from the tips. Here, the eigenvalue density decays slowly towards the tips from the maxima, as can be seen for $\rho\approx 0.76$ in Fig. 2(c).
@@
The leading eigenvalues in the spectrum control the stability of the dynamics. For $\rho=0$, the spectrum is given by the circular law [1, 48], and the leading eigenvalues are at the leading edge of the circle. In the limit of large $N$, the real part of many eigenvalues simultaneously exceeds $1$ at the onset of instability of the quiescent fixed point where $g_{\text{eff}}\approx 1^{+}$, leading to chaos  [1]. The same appears to be true for all $|\rho|<\rho_{\text{c}}$: the real part of many eigenvalues simultaneously exceeds $1$ and chaotic dynamics emerge with high probability. For $\rho\approx\rho_{\text{f}}$, where the leading edge of the spectrum is nearly flat (vertical), dynamics has maximal dimensionality (see Fig. 1(d)). At this value of $\rho$, the probability of observing chaotic dynamics when the origin becomes unstable is also maximal, as shown in Fig. 1(c) (see also Fig. S5 of SM.)
@@
As $\rho$ increases for $\rho>\rho_{\text{f}}$, the leading edge of the eigenvalue distribution becomes increasingly pointed and the magnitudes of the imaginary parts of the eigenvalues decrease, thus the fluctuations in the dynamics slow. (See insets of Figs. 2(a) and 2(b).) As $\rho$ decreases for $\rho<\rho_{\text{f}}$, the eigenvalues near the leading edge accumulate near the two imaginary points, leading to stable oscillations near the onset of instability of the quiescent fixed point. In both cases, for $|\rho|>\rho_{\text{c}}$, when there are cusps at the tips of the vertices, eigenvalues do not accumulate at the edge of the support. Instead, the eigenvalue density decays toward the tips of the cusps. Because of these properties, the number of unstable eigenvalues increases slowly with $g_{\text{eff}}$ after the onset, even in the limit of large $N$. (See inset of Fig. 2(c).) This slow accumulation of eigenvalues corresponds with the high probability of observing fixed points or oscillations in the dynamics. Thus, at $\rho=\pm\rho_{\text{c}}$, where there is a morphological change at the leading edge of the eigenvalue support, we conjecture that there are critical points where phase transitions in the dynamics at onset occur. This conclusion is consistent with what is known about networks with pairwise correlations ($\alpha=2$), where chaos occurs at onset for all $|\rho|<1$, but trajectories converge to equilibria at $\rho=\rho_{\text{c}}=1$ [22, 49].
@@
Although chaotic dynamics do not occur at the onset of instability of the origin for strong correlations where $|\rho|>\rho_{\text{c}}$, chaotic dynamics does emerge when $g_{\text{eff}}$ is large, e.g., when $|\rho|\approx 0.75$ chaotic behavior is still observed with high probability when $g_{\text{eff}}\gtrsim 3$ in Fig. 1(c). The value of $g_{\text{eff}}$ for which chaotic dynamics is first observed decreases as $|\rho|\rightarrow\rho_{\text{c}}$, presumably to $\sim 1^{+}$ in the limit of large $N$. However, simulations are delicate in this region because the transient time increases with both $g_{\text{eff}}$ and $N$, and it becomes difficult to distinguish between chaotic behavior, and fixed point or oscillatory dynamics with long transients.
@@
Numerical investigations of networks with higher order correlations, $\alpha=4,5\ \text{and}\ 6$ confirm the existence of phase transitions in the dynamics at the onset of instability of the fixed point at the origin and critical points at $\rho=\pm\rho_{\text{c}}$ consistent with the observation in the case $\alpha=3$. However, at values of $|\rho|\gg\rho_{\text{c}}$, deviations from the Hypotrochoid Law become increasingly large with $\alpha$.
@@
Dynamics in the region where chaotic behavior is observed are strongly influenced by high-order correlations. Figure 3 shows the average attractor dimensionality and Lyapunov exponents, $\lambda_{\text{LE}},$ as functions of correlation order, $\alpha$, at fixed values of $|\rho|\approx 0.18$ and $g_{\text{eff}}$, obtained by averaging over $50$ realizations of networks with $N=1600$. (See Fig. S6 of SM for the corresponding representative eigenvalue distributions.) Dimensionality and Lyapunov exponents behave similarly: For positive values of $\rho$, these two quantities decrease monotonically with $\alpha$. For negative values of $\rho$, both dimensionality and Lyapunov exponents grow as the spectrum of eigenvalues flattens at the leading edge of the spectrum on the real axis, reaching a peak when $\rho=\rho_{\text{f}}$. At the correlation value $\rho_{\text{f}}$ the spectrum at this leading edge is the flattest. This value depends on the correlation order, $\alpha$, and for third-order correlations the spectrum is flattest at $\rho=-.18$. This is consistent with the value of $\rho$ where we observed that dimensionality and Lyapunov exponents reach their maximum values, as shown in Fig. 3. Similarly, dimensionality as a function of correlation also peaks at $\rho\approx\rho_{\text{f}}$ as shown in Fig. 1(d).
@@
Thus, we conjecture the phase diagram in Fig. 4 for the dynamics at the onset of instability of the quiescent fixed point as a function of $\rho$ and $\alpha$. Lines of critical points (solid lines) separate the chaotic phase (red) from the phases where fixed points (blue) and stable oscillations (green) occur. The dimensionality of the dynamics and the Lyapunov exponents are maximal along a line in the middle of the chaotic region (dashed line).
Figure 4: Phase diagram of dynamics at the onset of the instability of the origin as a function of correlation strength $\rho$ and correlation order $\alpha$ in the limit of $N\rightarrow\infty$. Solid lines correspond to $\pm\rho_{\text{c}}$, where the eigenvalue support of the synaptic connectivity matrix develops cusps. The dashed line corresponds to $\rho_{\text{f}}$, where the leading edge of the support is nearly flat and the dimensionality of the dynamics is maximal.